A consumer’s utility function is given by 𝑈(𝑥1

,𝑥2

) = 2𝑥1𝑥2 + 3𝑥1

. Where 𝑥1

and 𝑥2

denote the number of items of two goods 𝐺1

and 𝐺2

that are bought. Each item costs $1 for 

𝐺1

and $2 for 𝐺2

. Use Lagrange multipliers to find the maximum value of U if the 

consumer’s income is $83. Estimate the new optimal utility if the consumer’s income rises 

by $1


1
Expert's answer
2021-03-08T07:20:11-0500

"u(X1X2)= 2X1X2+3X1"

where G1=$1, G2=$2 and B=$83


forming the budget constraint:

"B=G1X1+G2X2"

83=X1+2X2


"U=2X1X2+3X1+\lambdaλ(83-X1-2X2)"


"UX1=2X2+3-\lambda=0 ..................(i)"

"UX2=2X1-2\lambda=0 ..................(ii)"

"U\lambda=83- X1+2X2=0 ..............(iii)"


solving simultaneously:X1=43 ; X2=20 "\lambda =43"

substituting the values into the Lagrangian function


"U=2(43)(20)+3(43)+43(83-43-2(20))"

"U=2(860)+3(43)+43(0)"

"U=1720+129+0"

"U= 1849"

if the consumer's income increases by $1, the new budget becomes $84 and the new budget constraint becomes: "84=X"1 +2"X"2


"84-X"1 "-2" "X"2

"U=2X1X2+3X1+\lambda(84-X1-2X2)"

"UX1=2X1X2+3-\lambda=0 ...........(i)"

"UX2=2X1-2\lambda=0 ..................(ii)"

"U\lambda=84 - X1+2X2=0 ............(iii)"

Solving simultaneously: X1=43.5 ; X2=20.25 and "\lambda" = 43.5


substituting the values into the Lagrangian function gives the new utility as:

"U=2(43.5)(20.25)+3(43.5)+43.5(83-43.5-2(20.25))"

"U=2(880.75)+3(43.5)+43.5(0)"

"U=1761.75+130.5+0"

"U= 1892.25"


An increase of $1 in the income of the consumer will cause an approximate change of 43 as suggested by the lagrangian multiplier.